OpenStudy (anonymous):

If the average of six consecutive even nos. is 13, wht is the smallest of these nos.?

Parth (parthkohli):

What will be the six consecutive numbers in terms of x? Tell me?

Parth (parthkohli):

Let's say that the first one is x. What will be the rest of them?

OpenStudy (anonymous):

Hahaha i dont have an idea XD Sorry. Because honestly, this is just my review, i dont know anything about it. I just analyze the solution XD

OpenStudy (unklerhaukus):

The consecutive even numbers: \[x,x+2,x+4,x+6,x+8,x+10\] the average: \[\langle x\rangle=\frac{x+(x+2)+(x+4)+(x+6)+(x+8)+(x+10)}{6}=13\]

OpenStudy (unklerhaukus):

\[\frac{6x+30}{6}=13\]

OpenStudy (anonymous):

Thanks! XD

OpenStudy (unklerhaukus):

\[x+5=13\]

OpenStudy (anonymous):

8 XD

OpenStudy (unklerhaukus):

\[\frac{(8+10+12+14+16+18)}{6}=\frac{78}{6}=13\]